OpenStudy (anonymous):

Find the indefinite integral: Integral of (e^(-x))(6+(e^(-x)) dx

OpenStudy (anonymous):

Which part would be u?

OpenStudy (anonymous):

is it the 6+e^(-x) part?

OpenStudy (anonymous):

Yep!

OpenStudy (kinggeorge):

You hardly need a u-sub here. Just multiply out, and integrate each part separately. \[\int\limits e^{-x}(6+e^{-x})\;\;dx\]\[=\int\limits 6e^{-x} \;\;dx +\int\limits e^{-x}\cdot e^{-x}\;\;dx\]\[=\int\limits 6e^{-x} \;\;dx +\int\limits e^{-2x}\;\;dx\]

OpenStudy (anonymous):

ooohhh, see that's what I was kind of thinking, but the section is has got me so stuck on subbing u.. lol

OpenStudy (kinggeorge):

Well, \(u=6+e^{-x}\) would also work.

OpenStudy (anonymous):

so.. a general question. When is it good to substitute and when is it better to just distribute and solve?

sam (.sam.):

I would use u=-x du=-dx \[-\int\limits e^u \left(e^u+6\right) \, du\] then sub again h=e^u dh=e^u du \[-\int\limits (h+6) \, dh\] \[-\int\limits 6 \, dh-\int\limits h \, dh\] \[-\frac{h^2}{2}-6 h+c\] substitube back \[-\frac{1}{2} e^{-2 x} \left(12 e^x+1\right)+c\]

OpenStudy (anonymous):

My opinion is set priority for separate them out first! If you can't, think the next step substitution!

OpenStudy (anonymous):

@KingGeorge - so i could take the integral from the multiplied out format

OpenStudy (anonymous):

@Chlorophyll thanks, that's a good rule to go by to know which to use

OpenStudy (kinggeorge):

You could definitely do that. It would be my preferred way, but only because I don't like doing u-subs very much.

OpenStudy (anonymous):

thanks everyone! makes more sense now :)

OpenStudy (kinggeorge):

You're welcome.

OpenStudy (anonymous):

Pretty much like if you can't substitute, then next step is Integral by Part :)

OpenStudy (kinggeorge):

And if you can't do by by parts... Something has gone horribly horribly wrong.

OpenStudy (anonymous):

gotcha.. you all are such lifesavers!

OpenStudy (anonymous):

Then trig subs :P

OpenStudy (anonymous):

ewwwww no talk about trig! lol ;)

OpenStudy (kinggeorge):

I would say that if you have to do trig subs, something has already gone horribly horribly wrong.

OpenStudy (anonymous):

lol, noted @KingGeorge ! :)

OpenStudy (anonymous):

Yup, that's why I mean! However sb is so fond of trig, to the point apply it instead of U sub simple exponent var :P

OpenStudy (anonymous):

so the integral of e^-x would be -e^-x?

OpenStudy (anonymous):

*what :) Glad that majority feel the same =)

OpenStudy (anonymous):

Yep, careful with negative !

OpenStudy (anonymous):

and then e^(-2x) would turn into -e^(-2x)?

OpenStudy (anonymous):

natural log gets me mixed up sometimes, that's why i'm checking

OpenStudy (kinggeorge):

You've got to bring the 2 down as well. So you should get -2e^(-2x)

OpenStudy (anonymous):

oohhh ok right.. thank you

OpenStudy (anonymous):

e^u = u' e^u u = -2x --> u' = -2

OpenStudy (anonymous):

and my answer would be: 12(e^(-x))+(e^(-2x))

OpenStudy (anonymous):

oh, and +C at the end

OpenStudy (kinggeorge):

I'm getting something slightly different.

OpenStudy (kinggeorge):

If you divided by 2 it would be correct. Go over your work and make sure you've accounted for everything like that.

OpenStudy (anonymous):

ok

OpenStudy (kinggeorge):

Divided by -2 actually.

OpenStudy (anonymous):

I got -6(integrand) of e^(-x) + -2(integrand) of e^(-2x), then I combined the contstants.. maybe those aren't combinable?

OpenStudy (anonymous):